# ProtChatGPT: Towards Understanding Proteins with Large Language Models

### v1

@@ 3.1 Architecture
where $d_{k}$ represents the dimensionality of the keys, softmax is the softmax activation function. The refined tokens $\boldsymbol{T}^{\prime}$ of the self-attention block can be written as:
$\boldsymbol{T}^{\prime}=\mathbf{W}_{o}\times(w\cdot\mathbf{V}),$ (3)
where $\cdot$ represents the dot product, $\times$ means the matrix multiplication, and $\mathbf{W}_{o}$ is the projection matrix for output. Given the sequence features from ESM-1b as $\boldsymbol{E}_{seq}$, tokens then interact with $\boldsymbol{E}_{seq}$ through cross-attention layers, which can be formulated as:
[…]
where $\boldsymbol{T}^{\prime\prime}$ represents the refined tokens after cross-attention, $\mathbf{W}_{q}^{\prime}$, $\mathbf{W}_{k}^{\prime}$, $\mathbf{W}_{v}^{\prime}$ and $\mathbf{W}_{o}^{\prime}$ are a new set of learning transformation matrices. Additionally, benefiting from the text transformer, tokens can further interact with the textual descriptions through the same self-attention layers. Finally, the PLP-former produces the output $\boldsymbol{E}_{seq}$ after a linear feed-forward layer (Vaswani et al., 2017). Depending on the pre-training task, we implement distinct attention masking strategies (Li et al., 2023) within the self-attention block to regulate the token-text interaction. The training details of PLP-former are given in Appendix A. In this way, PLP-Former can effectively select the most useful information for the LLM while removing irrelevant protein information. This reduces the burden of the LLM to learn protein-language alignment, thus mitigating the catastrophic forgetting problem.
Note that we apply PLP only to the sequence embeddings since the reported protein structures are much less than sequences. For example, there are 182K experimentally-determined structures in Protein Data Bank (PDB) (Berman et al., 2000) while 47M protein sequences in Pfam (Mistry et al., 2021). Thus we only use selected sequence-structure pairs during the second-stage training.
Multi-Level Projection Adapter.
For the second protein-to-text generative learning stage, we further design a multi-level projection adapter to harvest the LLM’s generative language capability. The adapter takes the pre-aligned sequence embedding $\boldsymbol{E}_{seq}$ from PLP-former and structure embedding $\boldsymbol{E}_{str}$ from ESM-IF1 as inputs, and acts as an information bottleneck to the LLM decoder, such that its output protein representation can be interpreted by the LLM. In practice, we use two individual Fully-Connected (FC) layers to linearly project the output protein embeddings into the same dimension as the question embedding of the LLM. They function as soft protein prompts that condition the LLM on protein representation from 1D and 3D levels.
Large Language Models (LLMs).
Finally, the projected protein prompts are prepended to the question prompts (text embeddings of user questions) through concatenation. In implementation, we deploy the Vicuna-13b (Chiang et al., 2023) as our LLM decoder, which employs the Transformer decoder (Vaswani et al., 2017) to model the conditional generation probability $p_{\theta}(n_{i}|n_{<i})$ in the language model. Specifically, given the context vector of tokens $N_{i-1}$ (i.e., protein embeddings along with user questions), the generated tokens $n_{i}$ (i.e., answers) are computed as follows:
$h_{0}=N_{i-1}\mathbf{W}_{e}+\mathbf{W}_{p},N_{i-1}=(n_{1},n_{2},...,n_{i-1})$ (5)
$h_{l}=transformer(h_{l-1})(1\leq l\leq m)$
$p(n_{i}|N_{i-1})=softmax(h_{m}\mathbf{W}_{e}^{T})$
where $m$ refers to the number of layers, $W_{e}$ denotes the token embedding matrix, $W_{p}$ represents the positional embedding matrix, and $transformer$ is a decoder layer consisting of a multi-head self-attention block and a position-wise feedforward network. During the second training stage, the generated tokens $n_{i}$ are replaced with publicly available descriptions for protein-to-text generative learning, while remaining empty in the testing stage, expecting ProtChatGPT to generate descriptive answers for the given protein and corresponding questions.
3.2 Protein-Language Pre-training Datasets
@@
In order to train ProtChatGPT, dedicated protein-specific training datasets are indispensable for our proposed two-stage training strategy. Details and examples of our training datasets are given in Appendix B. During the first stage, we adopt ProtDescribe dataset (Xu et al., 2023) to train the PLP-Transformer for protein-description representation learning. ProtDescribe dataset contains 553,052 aligned pairs of protein sequences and textual property descriptions such as protein names, functions, families, subcellular locations, etc. The whole dataset is used for the first-stage pre-training. After the first stage, PLP-Transformer is able to acquire the protein 1D representation that is most pertinent to the textual description.
@@
Nevertheless, despite covering the protein sequence positions, ProtDescribe de facto simply relies on textual descriptions to provide a rough indication of protein structural and functional similarity. It might be more straightforward to directly utilize structural information, especially considering that ESM-IF is specifically designed for protein geometric structures. Considering this problem, we resort to the RCSB-PDB Protein Description Dataset (Guo et al., 2023) which comprises 143,508 structure-description aligned pairs of proteins. We further expand this dataset with 1D sequences, enabling the joint training with our PLP-former. For every protein taken into account, we compile its 1D residue sequences, 3D atomic coordinates, along with corresponding descriptions to train our multi-level adapter. We randomly select 1,000 pairs of protein for evaluation, and the rest part of the dataset is used for our second-stage pre-training.
3.3 Aligned Protein-Text Generation
To enhance the model training with protein-text pairs, we utilize a specialized token prompt, following the conversational format used in Vicuna-13b:
$\begin{array}[]{l}\mathbf{Q}{:}<Protein><ProteinPrompts></Protein>\\ <QuestionPrompts>\\ \mathbf{A}{:}<Descriptions>\end{array}$
[…]
4 Experiments
4.1 Implementation Details
For our training setup, we freeze both the sequence and structure protein encoders, as well as the LLM decoder, solely focusing on training the PLP-former (first stage) and the projection adapter (second stage). This two-stage strategy ensures that the pre-trained models retain their learned knowledge while fine-tuning the projection layer to align the protein embeddings with the LLM’s requirements. For the first stage, the PLP-former is initialized with the pre-trained weights of PubMedBERT (Gu et al., 2021), whereas the cross-attention layers and learnable tokens are randomly initialized. Following the same setting as PubMedBERT, we use 32 learnable tokens with a dimension of 768, which is the same as the hidden dimension of the PLP-Former. We pre-train the PLP-former on the ProtDescribe dataset for 20K epochs with a batch size of 64. We use the AdamW (Loshchilov & Hutter, 2017) optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.98, and a weight decay of 0.05. We use a cosine learning rate decay with a peak learning rate of 1e-4, a minimum learning rate of 8e-5, and a linear warm-up of 5K iterations. For the second stage, we freeze the PLP-former and LLM, and train the projection adapter for 1K epochs with a batch size of 128. The minimum learning rate for the second stage is set as 5e-5. To avoid out-of-memory issues, we randomly trim excessively long protein sequences and structures to 1000 amino acids. All experiments are performed on 4 NVIDIA A100 (80GB) GPUs, our model with ESM-1b and Vicuna-13b requires 5.5 days for the first stage and 2 days for the second stage.
4.2 Results
4.2.1 Qualitative Results
@@
We first show some example conversations between the user and ProtChatGPT to indicate the promising results on both protein understanding and design tasks. For a fair comparison, we use the protein that does not appear in the training set of ProtChatGPT. Figure 3 shows some conversation examples of our ProtChatGPT related to protein understanding and design, as well as customized protein design tasks. It can be observed that our system comprehends the meaning of the question well, providing accurate and logically consistent responses that not only identify the types and characteristics of proteins, but also extend to exploring their related application areas and potential values. Additionally, when presented with a protein, ProtChatGPT can provide explanations regarding questions associated with protein understanding. This implies that researchers can swiftly survey the structures, functions, and other relevant properties through ProtChatGPT. Coupled with this comprehension capability, further judgments on mutations and disease association can be made about specific proteins, leading to the potential of targeted protein design in healthcare.
Figure 3: Dialogue examples of ProtChatGPT for protein understanding and analysis (a, b) as well as customized protein design (c) tasks.
Figure 4: Case studies on homologous proteins (a, b) and mutually exclusive functions (c).
4.3 Case Study
@@
To further validate the utility of our method, we conduct case studies on two specific examples, including homologous proteins and mutually exclusive functions.
4.3.1 Homologous Proteins
@@
4.3.2 Mutually Exclusive Functions
@@
4.3.3 Quantitative Comparisons
To validate the effectiveness of our method, we further conduct several quantitative experiments. To fully showcase the capability of ProtChatGPT, we randomly select 1,000 protein sequence-structure-description pairs from the RCSB-PDB Protein Description Dataset dataset to serve as the test set. Note that these testing protein pairs are not used during training for a fair comparison. We employ seven commonly used metrics in the image captioning and NLP domains to test the performance of ProtChatGPT. Detailed descriptions of these metrics can be found in Appendix C.
We first devise two variants to validate our contribution. (1) w/o structure: remove the supplementary branch of the 3D structure encoder, relying solely on sequence information for learning. (2) w/o PLP-former: remove the PLP-former and directly aligned the 1D and 3D embeddings using the adapter. As a fair comparison, we give the LLM with the same user question as: “Describe this protein in short”. It can be observed from Table 1 that both the structural encoder and the PLP-former play indispensable roles in supplementation and alignment, respectively. Notably, in contrast to common metrics like BLEU and METEOR, SPICE and BERTScore pay more attention to deeper semantic information, rather than just lexical and syntactic alignment. Particularly, we replace the original Bert encoder with PubMedBERT (Gu et al., 2021), a biomedical description-specific encoder pre-trained on large-scale datasets. To some extent, this indicator can reflect the scientific validity of the generated responses in the biomedicine domain.
Table 1: Quantitative Comparisons on the proposed multi-level encoders and PLP-former. $\uparrow$ indicates that a higher value corresponds to better performance. The best performances are marked in bold.
VariantMetric BLEU-1 normal-$\uparrow$ BLEU-4 normal-$\uparrow$ ROUGE-L normal-$\uparrow$ METEOR normal-$\uparrow$ CIDEr normal-$\uparrow$ SPICE normal-$\uparrow$
@@ 4.3.3 Quantitative Comparisons
BERTScore normal-$\uparrow$
w/o structure 0.457 0.311 0.405 0.237 0.504 0.231 0.335
w/o PLP-former 0.581 0.352 0.463 0.270 0.572 0.276 0.421
ProtChatGPT 0.610 0.394 0.489 0.291 0.638 0.316 0.457
@@
Figure 5: Comparison of fine-tuning of PLP-former and LLM decoder during the second-stage training. We compute the SPICE and PubMed BERTScore for semantic evaluation.
@@
Furthermore, we modified the two-stage training strategy of PLP in an attempt to achieve better protein-specific dialogue capabilities. As shown in Figure 5, we independently fine-tune the LLM decoder and PLP-former in the second stage with the RCSB-PDB dataset. The performance on two high-level semantic metrics SPICE and PubMed BERTScore indicate that further fine-tuning of both LLMs and PLP-former enhances the performance. Considering the computational cost, we adopt the fastest two-stage separate training strategy.
4.4 Limitation and Discussion
ProtChatGPT leverages the capabilities of LLMs for protein-specialized conversations. However, it inherits LLM’s potential language hallucination. It is an indispensable concern especially when it relates to protein research and healthcare. Although the two case studies in Section 4.3 demonstrate that our ProtChatGPT can distinguish some complex cases (e.g., homologous proteins and mutually exclusive functions), these capabilities heavily depend on the presence of similar examples in our training dataset, especially the description part. Given an unknown protein, ProtChatGPT might produce certain descriptions that sound correct but lack proper scientific verification, possibly leading researchers astray. This issue might be alleviated by training the model with more high-quality, aligned protein-text pairs, or aligning with more advanced LLMs in the future. In this manner, rigorous data processing and selection strategies should be implemented to ensure the validity and reliability of the training data. Feedback from domain experts is also important to refine the model. Combined with reinforcement or continual learning techniques, ProtChatGPT could keep improving the quality of its responses. Furthermore, training only one projection adapter might not provide enough capacity to learn extensive protein-text alignment. This issue could be alleviated by designing a more powerful adapter to facilitate the interactions between sequence and structure embeddings, or utilizing other powerful protein encoders such as GearNet (Zhang et al., 2023) and ESM-2 (Lin et al., 2023). In future work, addressing these issues and refining ProtChatGPT is essential. With ongoing improvements and regular expert feedback, ProtChatGPT has the potential to become a trusted assistant in protein research, offering more valuable insights for further investigations.
5 Conclusion
@@
In this paper, we introduce ProtChatGPT, an AI-based protein chat system to implement ChatGPT-like functionalities for the protein research field. ProtChatGPT marks the initial effort at bootstrapping Protein-Language Pre-training (PLP) from pre-trained LLMs for various protein-language tasks. PLP sits at the intersection between protein and language, which effectively and efficiently enables ProtChatGPT to harvest the off-the-shelf large models from both protein and natural language communities. Experiments suggest that ProtChatGPT holds potential for application in protein understanding and design. We hope this work can facilitate protein research and further inspire other scientific disciplines.
Appendix A Details on Protein-Language Representation Learning

### v2

@@ 3.3.2. Protein-Text Generation
4. Experiments
4.1. Dataset
4.1.1. Protein-Language Pretraining Dataset
@@
In order to train our ProtChatGPT, dedicated protein-specific training datasets are indispensable for our proposed progressive training strategy. First, we adopt ProtDescribe dataset (Xu et al., 2023) to train the PLP-Transformer for protein-description representation learning, which contains 553,052 aligned pairs of protein sequences and textual descriptions such as protein names, functions, families, subcellular locations, etc.
4.1.2. Instruction Tuning Datasets
@@
Nevertheless, despite covering the protein sequence positions, ProtDescribe de facto simply relies on textual descriptions to provide a rough indication of protein structural and functional similarity. It might be more straightforward to directly utilize structural information, especially considering that NetSurfP and ESM-IF are specifically designed for high-level protein structures. Considering this problem, we resort to the RCSB-PDB dataset (Guo et al., 2023) which comprises 143,508 tertiary structure-description aligned pairs of proteins. We further extract the protein sequence and expand this dataset with sequence and secondary structure labels. All secondary structure labels are constructed using DSSP (Dictionary of Protein Secondary Structure, see Appendix C).
Moreover, considering fine-tuning LLMs on other domains typically requires many task-specific examples, we further collect the corresponding literature from the RCSB PDB website based on the protein entry ID, and then extract the PubMed abstracts as an additional external web corpus. Given the protein-related document, we utilize a recent method Bonito (Nayak et al., 2024) to convert unannotated text into corresponding Q&A-specific instruction pairs. In summary, for every protein taken into account, we compile its residue sequences, secondary structure label, tertiary atomic coordinates, along with corresponding 10 paired Q&A descriptions. We randomly select 1,000 pairs of protein for evaluation, and the rest part of the dataset is used for multi-level protein alignment (Section 3.2.2) and instruction tuning on LLMs (Section 3.3). More details and examples of our training datasets are given in Appendix F.
4.2. Implementation Details
In this paper, we use a progressive strategy to ensure that the pre-trained models retain their learned knowledge while fine-tuning them to align the protein embeddings with the LLMs. Firstly, we freeze all the multi-level protein encoders, solely focusing on training the PLP-former with the ProtDescribe dataset. Then the protein context gating module is trained with the RCSB-PDB dataset through contrastive learning. Finally, we freeze the LLM decoder and train the projection adapter with the generated instruction dataset. More implementation details are given in Appendix G.
4.3. Qualitative Comparisons
Figure 3. Several conversation examples. (a, b): ProtChatGPT can fully explore the intrinsic properties of proteins and accurately understand user queries, enabling protein understanding and analysis. (c): ProtChatGPT also has the potential to assist in drug development through pathogenicity analysis, diagnostic simulation, and protein design.
@@ 4.3. Qualitative Comparisons
We first show some example conversations between the user and ProtChatGPT to indicate the promising results on both protein understanding and design tasks. For a fair comparison, we use the protein that does not appear in the training set of ProtChatGPT. Figure 3 shows several conversation examples of our ProtChatGPT related to protein understanding and design, as well as customized protein design tasks. It can be observed that our system comprehends the meaning of the question well, providing accurate and logically consistent responses that not only identify the types and characteristics of proteins, but also extend to exploring their related application areas and potential values. Additionally, when presented with a protein, ProtChatGPT can provide explanations regarding questions associated with protein understanding. This implies that researchers can swiftly survey the structures, functions, and other relevant properties through ProtChatGPT. Coupled with this comprehension capability, further judgments on mutations and disease association can be made about specific proteins, leading to the potential of targeted protein design in healthcare.
Figure 4. Challenging case studies. Benefitting from contrastive-based feature alignment and domain-specific instruction tuning, our proposed ProtChatGPT is capable of handling complex scenarios such as (a, b) homologous proteins and (c) mutually exclusive functions.
4.4. Case Study
@@
To further validate the utility of our method, we conduct case studies on two specific challenging examples, including homologous proteins and mutually exclusive functions.
4.4.1. Homologous Proteins
@@
4.4.2. Mutually Exclusive Functions
@@
4.5. Quantitative Comparisons
4.5.1. Protein Q&A
First, we provide comparisons for the protein Q&A task on the PDB-QA dataset (Guo et al., 2023). We compared our LL A MA3-8B based method with ProteinChat (Guo et al., 2023) (Vicuna-13B based) and ProtT3 (Liu et al., 2024) (Galactica-1.3B based). We followed the same test setting as ProtT3 and selected the exact match as the evaluation metric. SP stands for structure/property, SI stands for supplementary information. We directly used the public checkpoints. As shown in Table 1, our method surpasses other baselines due to the combined encoding of all three levels of protein structures. In contrast, ProteinChat only uses tertiary structure while ProtT3 only uses protein sequence information, which ignores the cross-level complementarity and may lead to insufficient protein feature extraction.
Table 1. Protein Q&A task on PDB-QA dataset with GT and pseudo structures. Best performances are marked in bold.
Model String Number Overall
SP SI SP SI
ProteinChat 7.2 10.3 28.9 19.2 15.5
ProteinChat (pseudo label) 7.8 9.5 27.4 17.5 14.2
ProtT3 85.4 69.8 47.2 39.4 65.0
Ours 88.1 73.3 49.6 42.7 68.8
Ours (pseudo label) 87.6 72.8 49.2 41.9 67.7
Considering that users might only have the amino acid sequence, pseudo secondary and tertiary structures can be generated using structural prediction models such as AlphaFold2 (Jumper et al., 2021). To simulate real-world application scenarios, we further provide more experiments on a large evaluation benchmark for the Protein Q&A task to verify the robustness under insufficient input. The original experiment was conducted on the PDB-QA dataset, in which the ground truth tertiary structures are directly obtained from RCSB-PDB (Guo et al., 2023). With the same test set, here we use the protein sequence to predict the pseudo tertiary structure through AlphaFold2. We also use the pseudo-secondary structures obtained from NetSurfP (Klausen et al., 2019) for experiments. It can be observed that even when using pseudo-structure labels as test inputs, our model still performs well. Additionally, we also conducted validation experiments on previous methods (ProtT3 is not applicable since it only utilizes sequence information) using the same pseudo inputs, which also resulted in good performance. This further demonstrates the feasibility of using the AlphaFold dataset (Varadi et al., 2022) as pseudo labels for inference.
4.5.2. Protein-text Retrieval
We also provide experiments on the cross-modal protein-text retrieval task, as shown in Table 2. We use the ProteinKG25 (Zhang et al., 2022b) testset and select ProtST (Xu et al., 2023), ProteinCLAP (Liu et al., 2023) and ProtT3 (Liu et al., 2024) for comparisons. As the ProteinKG25 dataset only contains amino acid sequences as the protein query, this retrieval task essentially tests the performance of our PLP-former. For a fair comparison, we follow the same retrieval setting as ProtT3, which uses PTC to obtain the top-k ranked candidates and PTM for re-ranking (see Appendix E for PTC and PTM).
Table 2. Protein-text retrieval task on ProteinKG25 dataset.
Model Protein to Text Text to Protein
Acc R@20 Acc R@20
ProtST 5.5 41.6 5.8 43.4
ProteinCLAP 39.0 89.4 39.3 89.7
ProtT3 55.8 91.7 55.6 91.7
Ours 60.4 93.3 60.1 93.3
Note that different from ProtT3 which only uses protein sequences as input, our ProtChatGPT introduces multi-level alignment from primary to tertiary structures, which leads to better protein representation (see Table 1), and further enables protein cross-level structure retrieval through subsequent contrastive learning (see Appendix H).
Table 3. Quantitative comparisons on the proposed multi-level protein encoding, PCG module and PLP-former with frozen LLMs. $\uparrow$ indicates that a higher value corresponds to better performance.
Method BLEU-1 $\uparrow$ BLEU-4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ CIDEr $\uparrow$ SPICE $\uparrow$
@@ 4.5.2. Protein-text Retrieval
BERTScore $\uparrow$
w/o tertiary structure 0.461 0.316 0.411 0.247 0.515 0.240 0.348
w/o secondary structure & PCG 0.509 0.332 0.445 0.250 0.544 0.265 0.411
w/o PLP-former 0.588 0.356 0.469 0.278 0.582 0.287 0.431
w/o instruction tuning 0.604 0.373 0.478 0.288 0.595 0.299 0.447
ProtChatGPT (Vicuna-13B) 0.618 0.399 0.493 0.299 0.640 0.321 0.465
ProtChatGPT (LL A MA3-8B) 0.643 0.417 0.515 0.326 0.687 0.346 0.487
@@
Figure 5. Comparison of fine-tuning of PLP-former and LLM decoders during the instruction tuning stage. We compute the SPICE and PubMed BERTScore for semantic evaluation.
4.5.3. Ablation Study
To validate the effectiveness of our method, we further conduct several ablation studies. To fully showcase the capability of ProtChatGPT, we randomly select 1,000 protein protein-instruction pairs from our refined RCSB-PDB dataset to serve as the test set. Note that these testing protein pairs are not used during training for a fair comparison. We employ seven commonly used metrics in the image captioning and NLP domains to test the performance of ProtChatGPT. Detailed descriptions of these metrics can be found in Appendix I.
We first devise four variants to validate our contribution. (1) w/o tertiary structure: remove the branch of the tertiary structure encoding, relying solely on sequence and secondary structure embedding. (2) w/o secondary structure&PCG: remove the secondary structure encoder and the Protein Context Gating (PCG) module, directly use sequence and tertiary structure embeddings for contrastive learning. (3) w/o PLP-former: remove the PLP-former and directly use the sequence embedding from ESM-1b without the injection of protein descriptions. (4) w/o instruction tuning: using the original single descriptions from the RCSB-PDB dataset without generating multiple instruction pairs. It can be observed from Table 3 that all levels of protein structure information as well as the protein-language alignment play indispensable roles in supplementation and alignment, respectively. Notably, in contrast to common metrics like BLEU and METEOR, SPICE and BERTScore pay more attention to deeper semantic information, rather than just lexical and syntactic alignment. Particularly, we replace the original Bert encoder with PubMedBERT (Gu et al., 2021), a biomedical description-specific encoder pre-trained on large-scale datasets. To some extent, this indicator can reflect the scientific validity of the generated responses in the biomedicine domain.
@@ 4.5.3. Ablation Study
Furthermore, we modified the training strategy of PLP and LLMs in an attempt to achieve better protein-specific dialogue capabilities. As shown in Figure 5, we further independently fine-tune the PLP-former and LLM decoders (using 16-bit LoRA (Hu et al., 2021) for Vicuna-13B and 4-bit QLoRA (Dettmers et al., 2024) for LL A MA3-8B) with our refined dataset with instruction pairs. The performance on two high-level semantic metrics SPICE and PubMed BERTScore indicate that further fine-tuning of both LLMs and PLP-former enhances the performance. More results under different LLMs and fine-tuning settings are given in Appendix H.
5. Conclusion
@@
In this paper, we introduce ProtChatGPT, an AI-based protein chat system, to implement ChatGPT-like functionalities for the protein research field. ProtChatGPT marks the initial effort at bootstrapping multi-level Protein-Language Pretraining (PLP) from pre-trained LLMs for various protein-language tasks. PLP sits at the intersection between protein and language, which effectively and efficiently enables ProtChatGPT to harvest the off-the-shelf large models from both protein and natural language communities. Experiments suggest that ProtChatGPT holds potential for application in protein understanding and design. We hope this work can facilitate protein research and further inspire other scientific disciplines.
Appendix
Appendix A: The limitation and discussion of ProtChatGPT.
Appendix B: The broader impact of ProtChatGPT.
Appendix C: We provide a brief introduction on proteins, including protein structures, homologous Proteins, mutually Exclusive Functions, and Dictionary of Protein Secondary Structure (DSSP).
Appendix D: Introduction and related work on Vision-Language Pretraining (VLP).
Appendix E: Details on the proposed Protein-Language Representation Learning (PLP).
Appendix F: Details and examples of the training Datasets.
Appendix G: More implementation details.
Appendix H: Additional experiments on LLM fine-tuning and protein retrieval.
Appendix I: Details on evaluation metrics used in this paper.
Appendix A Limitation and Discussion
ProtChatGPT leverages the capabilities of LLMs for protein-specialized conversations. However, it inherits LLM’s potential language hallucination. It is an indispensable concern especially when it relates to protein research and healthcare. Although the two case studies in Section 4.4 demonstrate that our ProtChatGPT can distinguish some complex cases (e.g., homologous proteins and mutually exclusive functions), these capabilities heavily depend on the presence of similar examples in our training dataset, especially the description part. Given an unknown protein, ProtChatGPT might produce certain descriptions that sound correct but lack proper scientific verification, possibly leading researchers astray. This issue might be alleviated by training the model with more high-quality, aligned protein-text pairs, or aligning with more advanced LLMs in the future. In this manner, rigorous data processing and selection strategies should be implemented to ensure the validity and reliability of the training data. Feedback from domain experts is also important to refine the model. Combined with reinforcement or continual learning techniques, ProtChatGPT could keep improving the quality of its responses. Furthermore, training only one projection adapter might not provide enough capacity to learn extensive protein-text alignment. This issue could be alleviated by designing a more powerful adapter to facilitate the interactions between sequence and structure embeddings or utilizing other powerful protein encoders such as GearNet (Zhang et al., 2023) and ESM-2 (Lin et al., 2023). In future work, addressing these issues and refining ProtChatGPT is essential. With ongoing improvements and regular expert feedback, ProtChatGPT has the potential to become a trusted assistant in protein research, offering more valuable insights for further investigations.
Appendix B Broader Impact
ProtChatGPT extends across multiple dimensions of protein research and biotechnology. By facilitating dialogue-based interactions with a protein-specific model, ProtChatGPT could democratize access to complex protein analysis, enabling researchers, regardless of their computational expertise, to engage deeply with protein data. This could accelerate discovery processes in drug design, genetic research, and disease understanding by providing intuitive, conversational access to sophisticated protein analyses and hypothesis generation. Furthermore, ProtChatGPT could also serve educational purposes, aiding in the training of next-generation scientists by providing an interactive learning tool that offers immediate feedback and explanations about protein structures and functions. This might inspire innovative approaches in both academia and industry, potentially leading to novel therapeutic strategies and an enhanced understanding of biological mechanisms.
Appendix C Brief Introduction on Proteins
Proteins are large molecules formed by amino acids linked through peptide bonds and are essential components of cells and organisms. They perform a range of critical functions in the body, from providing structural support to regulating biochemical processes. Proteins are made up of hundreds or thousands of smaller units called amino acids, which are attached to one another in long chains. There are 20 different types of amino acids that can be combined to make a protein. The sequence of amino acids determines the unique 3-dimensional structure and the specific function of each protein.
C.1. Protein structures
Proteins have four levels of structure:
  • Primary Structure: The sequence of amino acids in a polypeptide chain.
  • Secondary Structure: Localized areas stabilized by hydrogen bonds, such as alpha-helices and beta-sheets.
  • Tertiary Structure: The overall 3-dimensional structure of a single protein molecule; the spatial relationship of the secondary structures to one another.
  • Quaternary Structure: The structure formed by several protein molecules (polypeptide chains), usually called protein subunits in this context, which function as a single protein complex.
C.2. Homologous Proteins
Homologous proteins are proteins that share a common evolutionary origin, indicated by their similarity in sequence or structure across different species. These proteins may retain similar functions, suggesting their conservation through evolutionary processes. Homologous proteins can be categorized into two types: orthologs, which are proteins in different species that evolved from a common ancestral gene and perform similar functions; and paralogs, which are proteins within the same species that arose from gene duplication and have evolved new functions. The study of homologous proteins helps in understanding protein function, evolution, and the genetic relationships between species.
C.3. Mutually Exclusive Functions
Mutually exclusive functions of proteins refer to situations where different protein isoforms, arising from the same gene through alternative splicing, perform distinct and non-overlapping roles within a cell or organism. Each isoform might be involved in different pathways or cellular processes, ensuring that their functions do not overlap, thus providing a mechanism for regulating biological activities and responses to environmental changes. This concept is crucial for understanding how genetic information can give rise to multiple functional outcomes, allowing a single gene to contribute to various aspects of cellular function and organismal development
C.4. Dictionary of Protein Secondary Structure (DSSP)
DSSP (Kabsch and Sander, 1983) is a standard method used to describe the secondary structure of proteins, which has become one of the most widely used tools for analyzing protein structures. DSSP analyzes the atomic coordinates of proteins to determine the secondary structure type of each amino acid residue, such as $\alpha$-helix, $\beta$-strand, turn, etc. It also provides additional information, such as hydrogen bonds between residues, the backbone torsion angles ($\phi$ and $\psi$), and solvent accessibility.
The output format of DSSP typically consists of a text file, with each line corresponding to an amino acid residue in the protein structure. Each line contains a series of labels describing the secondary structure and its local environment for that residue. Generally, the DSSP output format includes the following key information:
  • Residue Number: Indicates the position number of each amino acid residue in the protein.
  • Residue Type: Specifies the type of amino acid residue, such as ALA, ARG, ASP, etc.
  • Chain Identifier: Indicates the identifier for different chains in the protein, usually a letter.
  • Secondary Structure: Indicates the secondary structure type for the residue, such as H ($\alpha$-helix), E ($\beta$-sheet), T (turn), etc.
  • Hydrogen Bonds: Describes the hydrogen bonding interactions between the residue and its neighboring amino acid residues.
  • Phi ($\phi$) and Psi ($\psi$) angles: Represents the backbone $\phi$ and $\psi$ angles for the residue, used to describe the protein’s backbone conformation.
  • Solvent Accessibility: Describes the extent of exposure of the residue surface, i.e., the degree of contact between the residue and the aqueous solvent.
Each line in the DSSP output file contains a combination of these pieces of information to describe the features and environment of each residue in the protein structure.
Appendix D Vision-Language Pretraining (VLP)
Data collected from different modalities generally offer distinct perspectives, frequently synergizing to yield a comprehensive understanding, enhancing the overall comprehension of the data. Vision-Language Pretraining (VLP) aims to learn multimodal foundation models, showing improved performance on various vision-and-language tasks (Radford et al., 2021a). Existing VLP methods can be roughly divided into representation learning-based and generative learning-based. Representation learning-based methods (Radford et al., 2021b; Jia et al., 2021; Yao et al., 2021; Li et al., 2022b, 2021) usually consider the image-text pairs as multi-modal views of the same semantics, and perform contrastive or multi-view learning for the alignment between multiple modalities. Generative learning-based methods  (Li et al., 2019; Lu et al., 2019; Chen et al., 2020; Li et al., 2020; Zhang et al., 2021; Wang et al., 2022; Zeng et al., 2021; Bao et al., 2022) aim to reconstruct the corrupted text (image) with the assistance of visual (text) modality through MLM-like objectives. For example, SimVLM (Wang et al., 2021) introduces a single prefix language modeling (PrefixLM) objective for exploiting large-scale weak supervision in VLP. CoCa (Yu et al., 2022) further verifies the representation ability of autoregressive language modeling in the vision-language domain. In this paper, we consider protein as a specialized biological language that encodes and communicates biological information through its amino acid sequences and interactions. Inspired by existing VLP methods (Li et al., 2023; Zhu et al., 2023), we first propose a representation learning-based Protein-Language Pretraining (PLP) framework to understand protein sequences via natural language instructions, and then design a generative learning-based multi-level adapter to generate the question-related answers combined with extra structural embeddings.
Appendix E Details on Protein-Language Representation Learning